A cylindrical container has a base area of 100 meters squared and is 12 meters high. If the container is one-third filled with water, what is the volume of the water in the container?
Answer by nerdybill(2453)

A cylindrical container has a base area of 100 meters squared and is 12 meters high. If the container is one-third filled with water, what is the volume of the water in the container?
.
Total volume of cylinder = "area of base" * "height"
Total volume of cylinder = 100*12 = 1200 cubic meters
.
If water fills 1/3 of cylinder the volume of water is:
(1/3)(1200) = 400 cubic meters